An apple falls from an apple tree growing on a 20° slope. The apple hits the ground with an impact velocity of 16.2 m/s straight downward. What is the component of the apple's impact velocity parallel to the surface of the slope?


1
Expert's answer
2021-04-26T17:13:41-0400
"v_x=v\sin{20}\\v_x=16.2\sin{20}=5.54\frac{m}{s}"